Accelerator Update for June 23, 2008 - June 25, 2008
Monday June 23
The day shift began with Operations monitoring the Tevatron (TeV) store 6241 (108.45E30) and the Recycler (RR) stash (5.2E10), with the Antiproton Source stacking (7.4E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:42 AM, a Booster vacuum pump (B:IPRF23) tripped off and brought down BRF13 & 14. Everything reset.
At 8:48 AM, LRF4 tripped off. It reset.
FESS personnel reported at 9:06 AM that they were working on the air conditioner at TeV sector D4.
Pbar transferred all its antiprotons to the Recycler at 10:05 AM.
Operations prepared Pbar for access at 10:23 AM.
A water expert switched Pbar over to the plate and frame heat exchanger at 11:08 AM.
Pbar came out of access and resumed stacking at 11:13 AM.
The Recycler study ended at 1:04 PM.
At 1:56 PM, a TeV RF station (TRF2) tripped off. This caused a 6% drop in luminosity.
At 2:37 PM, TeV store 6241 was lost when TeV sectors C4, D1, and D0 quenched.
Operations notified FESS at 2:45 PM that this would be a good time to replace the TeV sector B0-B pump motor.
Operations reported at 2:59 PM that one group of experts were investigating a communication loss at C4 and other experts were investigating the TeV RF.
Operations prepared CDF and D0 for access at 3:09 PM.
Water experts reported at 3:13 PM that they were valving in the LCW heat exchanger at TeV sector B1.
RF experts began to work on TRF2 at 3:25 PM.
Operations reported at 3:59 PM that the Recycler had stashed antiprotons three times during the day shift.
At 4:22 PM, Operations reported that antiprotons were lost in the MI during a stash. This was due to a RF problem.
At 4:33 PM, Operations reported that beam would not make it through MI. Operators and experts investigated.
RF experts reported at 4:55 PM that the TRF2 work was done.
Operations reported at 5 PM that Pbar had resumed stacking. The MI RF problem was fixed by rebooting the low level and mid level RF systems.
Operations disabled the MTest beam for the day at 6 PM.
D0 and CDF were both out of access by 6:29 PM.
The B0 pump work was completed by 6:56 PM.
Operations put the TeV through a dry squeeze at 8:29 PM.
At 9:27 PM, Operations reported that BRF3 had tripped off over 30 times since 5 PM. Operators contacted an expert who suggested adjustments to the scale factor and to the anode ramp program. Operators also made a slight adjustment to BRF3.
At 9:31 PM, Operations reported that the output of the Linac had dropped by almost 1mA. Operators tuned.
Operations put the TeV through a wet squeeze at 9:41 PM.
At 10:03 PM, Operations received an alarm for a leaking Kautzky valve at TeV sector A4. Operators investigated and reported an open circuit in one of the sensor lines; the Kautzky valve was okay. The picture on the right is a diagram of the Kautzky valve.
Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.
Tuesday June 24
The midnight shift began with Operations monitoring the stash (256.3E10), with Pbar stacking (12.7E10), and with BNB taking beam.
A TeV expert completed a TeV orbit study at 12:01 AM.
At 12:17 AM, BRF3 tripped off. It reset.
Operations reported at 12:40 AM that they had put the TeV into shot setup earlier on the midnight shift.
Operations established store 6245 at 2:24 AM with an initial luminosity of 203.1E30.
At 3:21 AM, Operations reported that a MI Lambertson (LAM52) had tripped off. Operators investigated and then called in an expert.
A power supply expert reported at 3:48 AM that he had reset the filter cabinet breaker for LAM52.
At 5:11 AM BRF3 tripped off for the first of two times in an hour. It reset.
At 5:27 AM, Operations reported that the P1 transfer line beam permit had tripped off three times due to problems with a MI trim magnet power supply (V701) and then LAM52 tripped off again. The power supply expert was still here and he reset the breaker for LAM52.
At 5:38 AM, Operations reported some odd behavior from LAM52. Operators investigated and discovered that LAM52 and V701 got their references from C468 cards in MI crate $5B. There were also Recycler correctors located in this crate.
At 6:41 AM, the MI52 Septa had stopped working properly. The VME was not receiving T-clock events. A controls expert investigated and had everything back online by 7:19 AM.
Operations established beam to MTest at 7:28 AM.
At 7:29 AM, a controls expert reported that there seemed to be an intermittent VME T-clock failure at MI52. CAMAC experts were investigating further.
Operations reported at 8:02 AM that the Recycler had stashed four times during the midnight shift.
Recycler experts began a study period at 8:33 AM.
At 8:45 PM, Operations reported that the BNB beam permit was tripping off.
FESS personnel reported at 9:04 that they were working on the air-handling unit at MI-65.
Water experts reported at 9:28 AM that they were turning on the TeV sector E1 heat exchanger.
At 10:50 AM, Operations reported that they had two Booster West Anode watchdog trips within the last 10 minutes.
At 11:19 AM, the Booster West Anode watchdog tripped off for the first of six times in about 90 minutes.
At 1:15 PM, an RF expert had valved in an RF dummy load at A0; it contaminated the A1 LCW system. The expert isolated the dummy load and monitored its resistivity.
Operations reported at 3:56 PM that the Recycler had stashed five times during the day shift.
Operations reported at 4:01 PM that FESS had lowered the regulation temperature for the Pbar 95 LCW system. This caused a change in the bend fields and required adjustments.
Operations disabled the MTest beam at 6:14 PM.
Operations terminated store 6245 at 7:08 PM.
Water experts began work at 7:14 PM on the TeV sector E1 LCW system.
Operations put the TeV into shot setup at 7:19 PM.
Water expert reported that the TeV sector E1 LCW work had finished at 7:42 PM.
Two controls experts began work on the MI TCLK repeater problem at 8:51 PM.
Operations established store 6246 at 8:56 PM with an initial luminosity of 264.21E30.
At 10:17 PM, Operations reported a bad control card for corrector magnets in MI-50. Operators replaced the card.
At 11:08 PM, Operations reported a problem with a Recycler kicker (KPS4A) that was holding up transfers. Operations contacted an expert. While operators were troubleshooting the problem, the sequencer conducted a transfer that stashed 24mA of antiprotons into the Recycler. Operations contacted an expert again.
Operations reported at 11:52 PM that the Recycler had stashed three times during the evening shift.
Wednesday June 25
The midnight shift began with Operations monitoring store 6246 (155.01E30) and the stash (24.6E10), with Pbar stacking (77.8E10), and with BNB taking beam.
At 12:04 AM, CDF reported that they had problem with a low-level power supply that was making all the data for this store bad. (Later they reported that their data was good, but they would need an access to replace the power supply.)
Operators reported at 1:05 AM that they had inspected R:KPS4A and found nothing wrong. Operations consulted with two other experts.
At 1:33 AM, BRF3 tripped off for the first of many times over the rest of the midnight shift. It reset.
Operations reported at 4:07 AM that the MI kicker KPS4A had two jobs: to kick antiprotons out of the Recycler and into MI and to abort the Recycler beam. Experts told Operations that this kicker was controlled differently than other kickers and that the MI-40 TCLK repeater work had caused the kicker voltage to become unclamped due to a timing problem; the kicker was usually clamped at zero except for transfers to MI. Operators adjusted the kicker to zero and then set it to the normal antiproton extraction value.
Operations established beam to MTest at 7:06 AM.
Operations reported at 7:54 AM that the MI Lambertson (LAM52) had stopped reading back. Operations contacted controls experts.
Operations reported at 8:03 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store. CDF will be allowed a 2-hour access after store 6246 is terminated, which will occur around mid-afternoon.
Comments and Suggestions
|last modified 06/25/2008 email Fermilab|
|Security, Privacy, Legal|